OpenStudy (anonymous):

solve the equation 3m+5=8m-105

OpenStudy (anonymous):

the answer is easy... 22

OpenStudy (ajprincess):

Bring the terms containing m to one side and numbers to the other side and solve.

OpenStudy (anonymous):

would you like to see work? is so: 3m+5=8m-105 add 105 3m+110=8m minus 3m 110=5m divide by 5 5/110=5m/5 22=m

OpenStudy (anonymous):

best answer please?